# A variational discretization method for mean curvature flows by the Onsager principle

### v1

@@ 6.1.2 The evolution of a flower-shaped curve
$\vec{x}(\alpha)=((2-2^{\sin(5\alpha)})\cos\alpha,(2-2^{\sin(5\alpha)})\sin\alpha)^{T},\quad 0\leq\alpha\leq 2\pi.$ (95)
We discretize the curve uniformly in $\alpha$ by setting $\vec{x}_{i}=\vec{x}(\alpha_{i})$ with $\alpha_{i}=\frac{2\pi}{n}i$. The evolution of the curve is plotted in Figure 7, and the change of the discrete energy function $\mathcal{E}_{h}(t)$ is illustrated in Figure 8. We can see that the curve gradually changes to a circular shape and also shrinks in length. We can also see how the discrete energy function gradually decreases as the curve evolves under mean curvature flow.
@@ 6.2 Volume preserving mean curvature flow
In this subsection, we show numerical simulations for the volume preserving mean curvature flows. We choose the same initial curve as that in Section 6.1.2. We discretize the curve (95) uniformly in $\alpha$ by setting $\vec{x}_{i}=\vec{x}(\alpha_{i})$ with $\alpha_{i}=\frac{2\pi}{n}i$. The evolution of the curve is plotted in Figure 9. We can see that the flower-shaped curve slowly becomes convex and eventually evolves into a circular shape. Due to the constraint that the enclosed area remains constant, the shape of the curve will not change after it becomes a circle, i.e., it will reach a stationary state. The discrete energy function $\mathcal{E}_{h}(t)$ changes with time as shown in Figure 10. We can see that the discrete energy function decreases in early stage and then approaches to a constant when the system goes to the stationary state.
@@ 6.3 Wetting problems
$\tilde{\Gamma}(0):\vec{x}(\alpha)=(\cos\alpha,\sin\alpha)^{T}\quad(0\leq\alpha\leq\pi).$ (96)
Under the constraint that the area enclosed by $\tilde{\Gamma}(t)$ and the $x$-axis remains constant, when $\tilde{\Gamma}(t)$ reaches a stationary state under mean curvature flow, the energy no longer changes, i.e., the changing rate of the energy over time ~˙$\dot{\tilde{\mathcal{E}}}=0$. From the expression (58) of $\dot{\tilde{\mathcal{E}}}$, the left and right contact angles of the curve at stationary state are equal to the Young’s angle.
We discretize the curve uniformly in $\alpha$ by setting $\vec{x}_{i}=\vec{x}(\alpha_{i})$ with $\alpha_{i}=\frac{\pi}{n}i$ and take the Young’s angle $\theta_{Y}$ as $\frac{2\pi}{3}$ and $\frac{\pi}{3}$ respectively. The evolution of the curve is plotted in Figure 11 while the change of the discrete energy function ~$\tilde{\mathcal{E}}_{h}$ over time is shown in Figure 12. We can see that the droplet gradually change its shape from a semi-circle to a circular shape with stationary contact angles. From Figure 12 we can see that the discrete energy function initially decreases to a constant value.
@@
The exact solutions of wetting problem can be affected by lateral displacement of the curve. We thus define the error as the difference of the stationary-state energy between the discrete curve and the smooth curve, i.e.
$err:=|\tilde{\mathcal{E}}-\tilde{\mathcal{E}}_{h}|.$ (97)
The convergence order is calculated as in Equation (94). We conducted numerical simulations with different parameters to assess the accuracy and convergence of the method. For the case where Young’s angle is $\theta_{Y}=\frac{2\pi}{3}$, we use a time step size of $\Delta t=10^{-5}$ and varied the number of nodes as $n=5,10,20,40,80$. The simulation was performed for $T=4$ time steps. The resulting errors and convergence order are summarized in Table 2. Similarly, for the case where Young’s angle is $\theta_{Y}=\frac{\pi}{3}$, we used the same time step size and varied the number of nodes as before. The simulation was performed for $T=5$ time steps, and the corresponding errors and convergence order are presented in Table 3. We could see the method has second order convergent with respect to the spacial mesh size.
Table 2: $\theta_{Y}=\frac{2\pi}{3}$, error and convergence order of the numerical method in Section 5.2.
@@
err 0.0553 0.0107 0.0023 5.0217e-04 9.3273e-05
order - 2.3731 2.2035 2.2075 2.4286
Table 3: $\theta_{Y}=\frac{\pi}{3}$, error and convergence order of the numerical method in Section 5.2.
@@
err 0.0624 0.0122 0.0025 4.9335e-04 5.9605e-05
order - 2.3525 2.2802 2.3494 3.0491
@@ 7 Conclusions
In this paper, we develop a novel finite element method for mean curvature flows by using the Onsager principle as an approximation tool. The key feature of the method is that the discrete scheme preserves the energy dissipation relations of the Onsager principle. This is important for many applications. We apply the method to three different problems, namely the simple mean curvature flow, the volume preserving mean curvature flow and a wetting problem on substrate. Numerical examples show that the method works well for all the three problems.
There are several problems we need to study in the future. Firstly, we apply an explicit Euler scheme for the time discretization. The scheme is stable only the time step is small enough. We need consider implicit or semi-implicit scheme to improve the stability of the scheme. Another important issue is on the quality of meshes of the discrete geometric flow. To avoid mesh degeneracy and entanglement, we add a penalty term to the energy. The technique works well only when we need choose the penalty parameter and the time step properly. It is known that some other methods have properties to keep the mesh equally distributed in arc length. It is interesting to combine these techniques with our method. In addition, it is interesting to do numerical analysis of the method and also extend the method ot higher order approximations and other geometric flows.
8 Acknowledgement
The work was partially supported by NSFC 11971469 and 12371415. We also thank Professor Wei Jiang for help discussions.
[…]
@@ 8 Acknowledgement
In sections 3-5, we derived multiple systems of equations. For convenience of the reader, we will show the specific form of Equations (26) (32) (50) and (85) below.
Firstly, for Equation (26), the coefficient matrix $A$ and the right-hand term $G$ are computed as follows:
$\begin{gathered}A=\begin{pmatrix}A_{11}&A_{12}&&&A_{1n}\\ A_{21}&A_{22}&A_{23}\\ \quad&\ddots&\ddots&\ddots\\ &&A_{n-1,n-2}&A_{n-1,n-1}&A_{n-1,n}\\ A_{n,1}&&&A_{n,n-1}&A_{n,n}\end{pmatrix}.\end{gathered}$
@@
$A_{ii}=\frac{|\vec{x}_{i}(t)-\vec{x}_{i-1}(t)|+|\vec{x}_{i+1}(t)-\vec{x}_{i}(t)|}{3}I_{2},\quad i=1,...n-1,$ (98)
$A_{i,i+1}=A_{i+1,i}=\frac{|\vec{x}_{i+1}(t)-\vec{x}_{i}(t)|}{6}I_{2},\quad i=1,...n-1;$
@@
The right-hand term $G=(\vec{g}_{1}^{T},\vec{g}_{2}^{T},...,\vec{g}_{n}^{T})^{T}$ with
$\vec{g}_{i}=-\frac{\partial\mathcal{E}_{h}}{\partial\vec{x}_{i}}=\frac{\vec{x}_{i+1}(t)-\vec{x}_{i}(t)}{|\vec{x}_{i+1}(t)-\vec{x}_{i}(t)|}-\frac{\vec{x}_{i}(t)-\vec{x}_{i-1}(t)}{|\vec{x}_{i}(t)-\vec{x}_{i-1}(t)|},\quad i=1,...,n-1,$ (100)
@@
$\vec{g}_{n}=-\frac{\partial\mathcal{E}_{h}}{\partial\vec{x}_{n}}=\frac{\vec{x}_{1}(t)-\vec{x}_{n}(t)}{|\vec{x}_{1}(t)-\vec{x}_{n}(t)|}-\frac{\vec{x}_{n}(t)-\vec{x}_{n-1}(t)}{|\vec{x}_{n}(t)-\vec{x}_{n-1}(t)|}.$ (101)
Secondly, for Equation (32), the right-hand term $G^{\delta}=((\vec{g}_{1}^{\delta})^{T},(\vec{g}_{2}^{\delta})^{T},...,(\vec{g}_{n}^{\delta})^{T})^{T}$ is computed as follows:
$g_{i}^{\delta}=-\frac{\partial\mathcal{E}_{h}^{\delta}}{\partial\vec{x}_{i}}$ $=\frac{\vec{x}_{i+1}-\vec{x}_{i}}{|\vec{x}_{i+1}-\vec{x}_{i}|}-\frac{\vec{x}_{i}-\vec{x}_{i-1}}{|\vec{x}_{i}-\vec{x}_{i-1}|}$ (102)
@@
Next, for Equation (50), the coefficient matrix $\hat{A}$ and the right-hand term $\hat{G}$ are computed as follows:
$\begin{gathered}\hat{A}=\begin{pmatrix}A_{11}&A_{12}&&&&&&A_{1n}&\vec{b}_{1,n+1}\\ A_{21}&A_{22}&A_{23}\\ \quad&\ddots&\ddots&\ddots&&&&&\vdots\\ &&A_{i,i-1}&A_{i,i}&A_{i,i+1}\\ &&&A_{i+1,i}&A_{i+1,i+1}&A_{i+1,i+2}\\ &&&&\ddots&\ddots&\ddots&&\vdots\\ &&&&&A_{n-1,n-2}&A_{n-1,n-1}&A_{n-1,n}&\vec{b}_{n-1,n+1}\\ A_{n,1}&&&&&&A_{n,n-1}&A_{n,n}&\vec{b}_{n,n+1}\\ \vec{b}_{1,n+1}^{T}&&\ldots&&&\ldots&\vec{b}_{n-1,n+1}^{T}&\vec{b}_{n,n+1}^{T}&0\end{pmatrix},\end{gathered}$
@@
and $\hat{G}=(\vec{g}_{1}^{T},\vec{g}_{2}^{T},...,\vec{g}_{n}^{T},g_{n+1})^{T}$. Where the formulations of $A_{ij}(1\leq i,j\leq n)$ and $\vec{g}_{i}\ (1\leq i\leq n)$ are given in Equations (98), (99), (100) and (101), respectively. Inaddition, we have $g_{n+1}=0$. The elements in last row and last column of $\hat{A}$ are given by
$\vec{b}_{1,n+1}=\frac{P}{2}[\vec{x}_{2}(t)-\vec{x}_{n}(t)],$ (103)
$\vec{b}_{i,n+1}=\frac{P}{2}[\vec{x}_{i+1}(t)-\vec{x}_{i-1}(t)],\quad i=2,...,n;$
where $\vec{b}_{i,n+1}(1\leq i\leq n)$ are two dimensional column vectors.
@@
Finally, for Equation (85), the coefficient matrix $\tilde{A}$ and the right-hand term $\tilde{G}$ are computed as follows:
@@
$\begin{gathered}\tilde{A}=\begin{pmatrix}\tilde{A}_{11}&\tilde{A}_{12}&&&&&&&\tilde{\vec{b}}_{1,n+1}\\ \tilde{A}_{21}&\tilde{A}_{22}&\tilde{A}_{23}\\ \quad&\ddots&\ddots&\ddots&&&&&\vdots\\ &&\tilde{A}_{i,i-1}&\tilde{A}_{i,i}&\tilde{A}_{i,i+1}\\ &&&\tilde{A}_{i+1,i}&\tilde{A}_{i+1,i+1}&\tilde{A}_{i+1,i+2}\\ &&&&\ddots&\ddots&\ddots&&\vdots\\ &&&&&\tilde{A}_{n-1,n-2}&\tilde{A}_{n-1,n-1}&\tilde{A}_{n-1,n}&\tilde{\vec{b}}_{n-1,n+1}\\ &&&&&&\tilde{A}_{n,n-1}&\tilde{A}_{n,n}&\tilde{\vec{b}}_{n,n+1}\\ \tilde{\vec{b}}^{T}_{1,n+1}&&\ldots&&&\ldots&\tilde{\vec{b}}^{T}_{n-1,n+1}&\tilde{\vec{b}}^{T}_{n,n+1}&0\end{pmatrix}.\end{gathered}$
@@
Where $\tilde{A}_{ij}(1\leq i,j\leq n)$ are square matrixes of order 2, and $\tilde{\vec{b}}_{i,n+1}(1\leq i\leq n)$ are two dimensional column vectors. The diagonal elements of the $2n$-order principal subblock of $\tilde{A}$ are computed as
$\tilde{A}_{11}=\left(\begin{array}[]{cc}\frac{\xi_{0}x_{2}^{(2)}(t)\sin\theta_{0}}{3}+\xi_{1}&0\\ 0&1\end{array}\right)\quad for\quad i=1$ (104)
$\tilde{A}_{ii}=\xi_{0}\frac{|\vec{x}_{i}(t)-\vec{x}_{i-1}(t)|+|\vec{x}_{i+1}(t)-\vec{x}_{i}(t)|}{3}I_{2}\quad for\quad 2\leq i\leq n-1$
$\tilde{A}_{nn}=\left(\begin{array}[]{cc}\frac{\xi_{0}x_{n-1}^{(2)}(t)\sin\theta_{L}}{3}+\xi_{1}&0\\ 0&1\end{array}\right)\quad for\quad i=n.$
where
$\sin\theta_{0}=\frac{x_{2}^{(2)}(t)}{|\vec{x}_{2}(t)-\vec{x}_{1}(t)|},\quad\sin\theta_{L}=\frac{x_{n-1}^{(2)}(t)}{|\vec{x}_{n}(t)-\vec{x}_{n-1}(t)|}.$ (105)
The minor diagonal elements of the $2n$-order principal subblock of $\tilde{A}$ are computed as
$\tilde{A}_{21}=\tilde{A}_{12}^{T}=(\frac{\xi_{0}x_{2}^{(2)}(t)\vec{n}_{1}}{6},\vec{0})\quad for\quad i=1$ (106)
$\tilde{A}_{i,i+1}=\tilde{A}_{i+1,i}^{T}=\xi_{0}\frac{|\vec{x}_{i+1}(t)-\vec{x}_{i}(t)|}{6}I_{2}\quad for\quad 2\leq i\leq n-2$
$\tilde{A}_{n-1,n}=\tilde{A}_{n,n-1}^{T}=(-\frac{\xi_{0}x_{n-1}^{(2)}(t)\vec{n}_{n-1}}{6},\vec{0})\quad for\quad i=n-1.$
The last row and column of $\tilde{A}$ are calculated as
$\tilde{\vec{b}}_{1,n+1}=(\frac{x_{2}^{(2)}(t)}{2},0)^{T}\quad for\quad i=1$ (107)
$\tilde{\vec{b}}_{i,n+1}=\frac{P}{2}[\vec{x}_{i+1}(t)-\vec{x}_{i-1}(t)]\quad for\quad 2\leq i\leq n-1.$
$\tilde{\vec{b}}_{n,n+1}=(-\frac{x_{n-1}^{(2)}(t)}{2},0)^{T}\quad for\quad i=n.$

### v3

@@ 6.1.2 The evolution of a flower-shaped curve
$\vec{x}(\alpha)=(2-2^{\sin(5\alpha)})(\cos\alpha,\sin\alpha)^{\top},\quad 0\leq\alpha\leq 2\pi.$ (46)
We discretize the curve uniformly in $\alpha$ by setting $\vec{x}_{i}=\vec{x}(\alpha_{i})$ with $\alpha_{i}=\frac{2\pi}{n}i$. The evolution of the curve is plotted in Fig. 7, and the change of the discrete energy function $\mathcal{E}_{h}(t)$ is illustrated in Fig. 8. We can see that the curve gradually changes to a circular shape and also shrinks in length. We can also see how the discrete energy function gradually decreases as the curve evolves under mean curvature flow.
@@
6.1.3 Effect of the penalty term
In section 3.3, we have added a penalty term into the discrete energy function $\mathcal{E}_{h}$ to avoid mesh degeneracy. Here, we will show the effect of the penalty. Define a mesh ratio indicator (MRI) $\Psi(t)$ of the discrete curve $\Gamma_{h}(t)$ (as in [53]) that
$\Psi(t)=\frac{\max\limits_{1\leq i\leq n}|\vec{x}_{i+1}-\vec{x}_{i}|}{\min\limits_{1\leq i\leq n}|\vec{x}_{i+1}-\vec{x}_{i}|}.$
For the circular curve case, the change of the MRI function $\Psi(t)$ is shown in Fig. 9. Notably, the MRI function $\Psi(t)$ remains constant $1$ (i.e. the mesh maintain a uniform distribution) for both situations with or without the penalty term. In this case, the penalty term is not necessary. However, for the flower-shaped curve case, the situation is different, as shown in Fig. 10. It is observed that without the penalty term, $\Psi(t)$ changes drastically over time, indicating very poor mesh property. In comparison, if we include the penalty term, the change of $\Psi(t)$ is mild and goes to $1$ gradually. Therefore, adding a penalty term in the energy significantly improves the mesh quality during the evolution of a flower-shaped curve under mean curvature flow. We also test the effect of the the penalty term for the numerical experiments in next subsections. The observations are similar. The penalty term seems not necessary for smooth curves like circles while it helps to improve the mesh quality for more singular curves.
(a) $\mathcal{E}_{h}(t)$ does not include the penalty term.
(b) Adding the penalty term to $\mathcal{E}_{h}(t)$.
Figure 9: The change of the MRI function $\Psi(t)$ over time for the circular curve.
(a) $\mathcal{E}_{h}(t)$ does not include the penalty term.
(b) Adding the penalty term to $\mathcal{E}_{h}(t)$.
Figure 10: The change of the MRI function $\Psi(t)$ over time for the flower-shaped curve.
@@ 6.2 Volume preserving mean curvature flow
In this subsection, we show numerical simulations for the volume preserving mean curvature flow. We choose the same initial curve as that in Section 6.1.2. We discretize the curve (46) uniformly in $\alpha$ by setting $\vec{x}_{i}=\vec{x}(\alpha_{i})$ with $\alpha_{i}=\frac{2\pi}{n}i$. The evolution of the curve is plotted in Fig. 11. We can see that the flower-shaped curve slowly becomes convex and eventually evolves into a circular shape. Due to the constraint that the enclosed area remains constant, the shape of the curve will not change after it becomes a circle, i.e., it will reach a stationary state. The discrete energy function $\mathcal{E}_{h}(t)$ changes with time as shown in Fig. 12. We can see that the discrete energy function decreases in early stage and then approaches to a constant when the system goes to the stationary state.
@@ 6.3 Wetting problems
$\tilde{\Gamma}(0):\vec{x}(\alpha)=(\cos\alpha,\sin\alpha)^{\top},\quad 0\leq\alpha\leq\pi.$ (47)
Under the constraint that the area enclosed by $\tilde{\Gamma}(t)$ and the $x$-axis remains constant, when $\tilde{\Gamma}(t)$ reaches a stationary state under mean curvature flow, the energy no longer changes. From the expression (30d), the left and right contact angles of the curve at stationary state are equal to the Young’s angle.
We discretize the curve uniformly in $\alpha$ by setting $\vec{x}_{i}=\vec{x}(\alpha_{i})$ with $\alpha_{i}=\frac{\pi}{n}i$ and take the Young’s angle $\theta_{Y}$ as $\frac{2\pi}{3}$ and $\frac{\pi}{3}$ respectively. The evolution of the curve is plotted in Fig. 13 while the change of the discrete energy function ~$\tilde{\mathcal{E}}_{h}$ (with $C=0$ in Eq. (35)) over time is shown in Fig. 14. We can see that the droplet gradually change its shape from a semi-circle to a circular shape with stationary contact angles. From Fig. 14 we can see that the discrete energy function initially decreases to a constant value.
@@
The exact solutions of wetting problems can be affected by lateral displacement of the curve. We thus define the error as the difference of the stationary-state energy between the discrete curve and the smooth curve, i.e.
$err_{2}:=|\tilde{\mathcal{E}}-\tilde{\mathcal{E}}_{h}|.$ (48)
The convergence order is calculated as in Eq. (45). We conducted numerical simulations with different parameters to evaluate the accuracy and convergence of the method, setting $\xi_{0}=\xi_{1}=1$. For the case where Young’s angle is $\theta_{Y}=\frac{2\pi}{3}$, we use a time step size of $\Delta t=10^{-5}$ and varied the number of nodes as $n=5,10,20,40,80$. The simulation was performed until $T=4$. The resulting errors and convergence order are summarized in Table 2. Similarly, for the case where Young’s angle is $\theta_{Y}=\frac{\pi}{3}$, we used the same time step size and varied the number of nodes as before. The simulation was performed until $T=5$, and the corresponding errors and convergence order are presented in Table 3. We could see the method has second order convergent with respect to the spacial mesh size.
Table 2: The errors on the energy and the convergence order for $\theta_{Y}=\frac{2\pi}{3}$.
@@
$err_{2}$ 5.5339e-02 1.0682e-02 2.3179e-03 5.0081e-04 9.2443e-05
$Order$ 2.37 2.20 2.21 2.44
Table 3: The errors on the energy and the convergence order for $\theta_{Y}=\frac{\pi}{3}$.
@@
$err_{2}$ 6.2309e-02 1.2166e-02 2.4918e-03 4.8600e-04 5.6509e-05
$Order$ 2.36 2.29 2.36 3.10
Notice that the energy is a very weak measure on the accuracy of the numerical solution. We could also consider the convergence of the error under some stronger measures, like the manifold distance [53], i.e.
$err_{3}:=|(\Omega_{h}\backslash\Omega)\cup(\Omega\backslash\Omega_{h})|=|\Omega_{h}|+|\Omega|-2|(\Omega_{h}\cap\Omega)|.$ (49)
Here $\Omega_{h}$ is the region enclosed by the discrete curve $\tilde{\Gamma}_{h}$ and the horizontal line, and $\Omega$ is the region enclosed by the smooth curve $\tilde{\Gamma}$ and the horizontal line. However, we find that the convergence rate is not optimal for the error. The main reason is that we are comparing with the analytical solution in stationary state, which is obtained by the numerical solution at final time $T$ even the energy seems unchanged anymore. To accelerate the process of the droplet going to stationary state, we decrease the friction coefficients $\xi_{0}$ and $\xi_{1}$. For the case $\theta_{Y}=\frac{2\pi}{3}$, we set $\xi_{0}=\xi_{1}=0.1$, with a time step size of $\Delta t=5\times 10^{-6}$. For the case $\theta_{Y}=\frac{\pi}{3}$, we set $\xi_{0}=\xi_{1}=0.02$, with a time step size of $\Delta t=10^{-6}$. The end time is the same as the previous tests. The numerical errors and convergence order are given in Tables 4 and 5. We can see that our method gives optimal convergence order with respect to spatial mesh size for the errors measured in manifold distance.
Table 4: The errors in manifold distance and the convergence order for $\theta_{Y}=\frac{2\pi}{3}$.
n 5 10 20 40 80
$err_{3}$ 1.7023e-01 3.4356e-02 7.6659e-03 1.7409e-03 4.4573e-04
$Order$ 2.31 2.16 2.14 1.97
Table 5: The errors in manifold distance and the convergence order for $\theta_{Y}=\frac{\pi}{3}$.
n 5 10 20 40 80
$err_{3}$ 1.5658e-01 3.1706e-02 7.1495e-03 1.6998e-03 4.1545e-04
$Order$ 2.30 2.15 2.07 2.03
@@ 7 Conclusions
In this paper, we develop a novel finite element method for mean curvature flows by using the Onsager principle as an approximation tool. The key feature of the method is that the semi-discrete scheme preserves the energy dissipation relations of the Onsager principle. This is important for many applications. We apply the method to three different problems, namely the simple mean curvature flow, the volume preserving mean curvature flow and a wetting problem on substrate. Numerical examples show that the method works well for all the three problems.
There are several problems we need to study in the future. Firstly, we apply an explicit Euler scheme for the time discretization. The scheme is stable only the time step is small enough. We need consider implicit or semi-implicit scheme to improve the stability of the scheme. Another important issue is on the quality of meshes of the discrete geometric flow. To avoid mesh degeneracy and entanglement, we add a penalty term to the energy. The technique works well only when we need choose the penalty parameter and the time step properly. It is known that some other methods have properties to keep the mesh equally distributed in arc length. It is interesting to combine these techniques with our method. In addition, it is interesting to do numerical analysis of the method and also extend the method to higher order approximations and other geometric flows.
Acknowledgments
The work was partially supported by NSFC 11971469 and 12371415. We would also like to acknowledge the support of the Beijing Natural Science Foundation (Grant No. Z240001). Additionally, we thank Professor Wei Jiang for his helpful discussions.
@@ Acknowledgments
In sections 3-5, we derived multiple systems of equations. For convenience of the reader, we will show the specific form of Eqs. (12), (15), (24) and (38) below.
Firstly, for Eq. (12), the coefficient matrix $A$ and the right-hand term $G$ are computed as follows:
$\begin{gathered}A=\begin{pmatrix}A_{11}&A_{12}&&&A_{1n}\\ A_{21}&A_{22}&A_{23}\\ \quad&\ddots&\ddots&\ddots\\ &&A_{n-1,n-2}&A_{n-1,n-1}&A_{n-1,n}\\ A_{n,1}&&&A_{n,n-1}&A_{n,n}\end{pmatrix},\end{gathered}$
@@
$A_{ii}=\frac{|\vec{x}_{i}(t)-\vec{x}_{i-1}(t)|+|\vec{x}_{i+1}(t)-\vec{x}_{i}(t)|}{3}I_{2},$ $1\leq i\leq n-1,$ (A.1)
$A_{i,i+1}=A_{i+1,i}=\frac{|\vec{x}_{i+1}(t)-\vec{x}_{i}(t)|}{6}I_{2},$ $1\leq i\leq n-1,$
@@
The right-hand term $G=(\vec{g}_{1}^{\top},\vec{g}_{2}^{\top},...,\vec{g}_{n}^{\top})^{\top}$ with
$\vec{g}_{i}=-\frac{\partial\mathcal{E}_{h}}{\partial\vec{x}_{i}}=\frac{\vec{x}_{i+1}(t)-\vec{x}_{i}(t)}{|\vec{x}_{i+1}(t)-\vec{x}_{i}(t)|}-\frac{\vec{x}_{i}(t)-\vec{x}_{i-1}(t)}{|\vec{x}_{i}(t)-\vec{x}_{i-1}(t)|},\quad 1\leq i\leq n-1,$ (A.3)
@@
$\vec{g}_{n}=-\frac{\partial\mathcal{E}_{h}}{\partial\vec{x}_{n}}=\frac{\vec{x}_{1}(t)-\vec{x}_{n}(t)}{|\vec{x}_{1}(t)-\vec{x}_{n}(t)|}-\frac{\vec{x}_{n}(t)-\vec{x}_{n-1}(t)}{|\vec{x}_{n}(t)-\vec{x}_{n-1}(t)|}.$ (A.4)
By applying mass lumping to Eq. (12), we obtain the classical semi-discrete Dziuk scheme.
Secondly, for Eq. (15), the right-hand term $G^{\delta}=((\vec{g}_{1}^{\delta})^{\top},(\vec{g}_{2}^{\delta})^{\top},...,(\vec{g}_{n}^{\delta})^{\top})^{\top}$ is computed as follows:
$g_{i}^{\delta}=-\frac{\partial\mathcal{E}_{h}^{\delta}}{\partial\vec{x}_{i}}$ $=\frac{\vec{x}_{i+1}-\vec{x}_{i}}{|\vec{x}_{i+1}-\vec{x}_{i}|}-\frac{\vec{x}_{i}-\vec{x}_{i-1}}{|\vec{x}_{i}-\vec{x}_{i-1}|}$
@@
Next, for Eq. (24), the coefficient matrix $\hat{A}$ and the right-hand term $\hat{G}$ are computed as follows:
$\begin{gathered}\hat{A}=\begin{pmatrix}A_{11}&A_{12}&&&&&&A_{1n}&\vec{b}_{1,n+1}\\ A_{21}&A_{22}&A_{23}\\ \quad&\ddots&\ddots&\ddots&&&&&\vdots\\ &&A_{i,i-1}&A_{i,i}&A_{i,i+1}\\ &&&A_{i+1,i}&A_{i+1,i+1}&A_{i+1,i+2}\\ &&&&\ddots&\ddots&\ddots&&\vdots\\ &&&&&A_{n-1,n-2}&A_{n-1,n-1}&A_{n-1,n}&\vec{b}_{n-1,n+1}\\ A_{n,1}&&&&&&A_{n,n-1}&A_{n,n}&\vec{b}_{n,n+1}\\ \vec{b}_{1,n+1}^{\top}&&\ldots&&&\ldots&\vec{b}_{n-1,n+1}^{\top}&\vec{b}_{n,n+1}^{\top}&0\end{pmatrix},\end{gathered}$
@@
and $\hat{G}=(\vec{g}_{1}^{\top},\vec{g}_{2}^{\top},...,\vec{g}_{n}^{\top},g_{n+1})^{\top}$. Where the formulations of $A_{ij}\ (1\leq i,j\leq n)$ and $\vec{g}_{i}\ (1\leq i\leq n)$ are given in Eqs. (A.1), (A.2), and (A.3), (A.4), respectively. In addition, we have $g_{n+1}=0$. The elements in last column of $\hat{A}$ are given by
$\vec{b}_{i,n+1}=\frac{P}{2}[\vec{x}_{i+1}(t)-\vec{x}_{i-1}(t)],\quad for\quad 2\leq i\leq n-1,$
$\vec{b}_{1,n+1}=\frac{P}{2}[\vec{x}_{2}(t)-\vec{x}_{n}(t)],\quad\vec{b}_{n,n+1}=\frac{P}{2}[\vec{x}_{1}(t)-\vec{x}_{n-1}(t)].$
@@
Finally, for Eq. (38), the coefficient matrix $\tilde{A}$ and the right-hand term $\tilde{G}$ are computed as follows:
@@
$\begin{gathered}\tilde{A}=\begin{pmatrix}\tilde{A}_{11}&\tilde{A}_{12}&&&&&&&\tilde{\vec{b}}_{1,n+1}\\ \tilde{A}_{21}&\tilde{A}_{22}&\tilde{A}_{23}\\ \quad&\ddots&\ddots&\ddots&&&&&\vdots\\ &&\tilde{A}_{i,i-1}&\tilde{A}_{i,i}&\tilde{A}_{i,i+1}\\ &&&\tilde{A}_{i+1,i}&\tilde{A}_{i+1,i+1}&\tilde{A}_{i+1,i+2}\\ &&&&\ddots&\ddots&\ddots&&\vdots\\ &&&&&\tilde{A}_{n-1,n-2}&\tilde{A}_{n-1,n-1}&\tilde{A}_{n-1,n}&\tilde{\vec{b}}_{n-1,n+1}\\ &&&&&&\tilde{A}_{n,n-1}&\tilde{A}_{n,n}&\tilde{\vec{b}}_{n,n+1}\\ \tilde{\vec{b}}^{\top}_{1,n+1}&&\ldots&&&\ldots&\tilde{\vec{b}}^{\top}_{n-1,n+1}&\tilde{\vec{b}}^{\top}_{n,n+1}&0\end{pmatrix},\end{gathered}$
@@
where $\tilde{A}_{ij}\ (1\leq i,j\leq n)$ are square matrices of order 2, and $\tilde{\vec{b}}_{i,n+1}\ (1\leq i\leq n)$ are two dimensional column vectors. The subblock on the diagonal of $\tilde{A}$ are computed as
$\tilde{A}_{ii}=\xi_{0}\frac{|\vec{x}_{i}(t)-\vec{x}_{i-1}(t)|+|\vec{x}_{i+1}(t)-\vec{x}_{i}(t)|}{3}I_{2},\quad 2\leq i\leq n-1,$
$\tilde{A}_{11}=\left(\begin{array}[]{cc}\xi_{0}\frac{|\vec{x}_{2}(t)-\vec{x}_{1}(t)|}{3}+\xi_{1}&0\\ 0&1\end{array}\right),\quad\tilde{A}_{nn}=\left(\begin{array}[]{cc}\xi_{0}\frac{|\vec{x}_{n}(t)-\vec{x}_{n-1}(t)|}{3}+\xi_{1}&0\\ 0&1\end{array}\right).$
The subblocks on the superdiagonal and subdiagonal of $\tilde{A}$ are computed as
$\tilde{A}_{i,i+1}=\tilde{A}_{i+1,i}^{\top}=\xi_{0}\frac{|\vec{x}_{i+1}(t)-\vec{x}_{i}(t)|}{6}I_{2},\quad 2\leq i\leq n-2,$
$\tilde{A}_{21}=\tilde{A}_{12}^{\top}=\left(\begin{array}[]{cc}\xi_{0}\frac{|\vec{x}_{2}(t)-\vec{x}_{1}(t)|}{6}&0\\ 0&0\end{array}\right),\quad\tilde{A}_{n-1,n}=\tilde{A}_{n,n-1}^{\top}=\left(\begin{array}[]{cc}\xi_{0}\frac{|\vec{x}_{n}(t)-\vec{x}_{n-1}(t)|}{6}&0\\ 0&0\end{array}\right).$
The last column of $\tilde{A}$ are calculated as
$\tilde{\vec{b}}_{i,n+1}=\frac{P}{2}[\vec{x}_{i+1}(t)-\vec{x}_{i-1}(t)],\quad 2\leq i\leq n-1,$
$\tilde{\vec{b}}_{1,n+1}=(\frac{x_{2}^{(2)}(t)}{2},0)^{\top},\quad\tilde{\vec{b}}_{n,n+1}=(-\frac{x_{n-1}^{(2)}(t)}{2},0)^{\top}.$